OpenStudy (anonymous):

Given that a = 10, b = 2x + 5, and ab=6xsquared+11x-10, find x.

OpenStudy (anonymous):

\[ab = 6x^2 + 11x - 10\]

OpenStudy (anonymous):

http://www.tiger-algebra.com/drill/6x~2-11x-10=0/ Don't know if it's much help.

OpenStudy (anonymous):

What does x equal?

OpenStudy (anonymous):

\((10)(2x + 5) = 6x^2 + 11x - 10\)

OpenStudy (anonymous):

A. 8/3 B. 7/2 C. 4 D. 6

OpenStudy (anonymous):

\(\ 20x + 50 = 6x^2 + 11x - 10\)

OpenStudy (anonymous):

\(\ 20x = 6x^2 + 11x - 60\)

OpenStudy (anonymous):

\(\ 6x^2 - 9x - 60\)

OpenStudy (anonymous):

Graph \(\ 6x^2 - 9x - 60\) here: https://www.desmos.com/calculator

OpenStudy (anonymous):

Look at the x-intercepts and those will be your answers.

OpenStudy (anonymous):

@ashlyn000

OpenStudy (anonymous):

@ashlyn000 You there?

OpenStudy (anonymous):

@;Ron Has it wrong..

OpenStudy (anonymous):

Lol I tryed :P

OpenStudy (anonymous):

He inputted '\(\ 6x^2 + 11x - 10 = 0\)' which is wrong. You're supposed to input '\(\ (10)(2x + 5) = 6x^2 + 11x - 10\)'

OpenStudy (anonymous):

The answers to \(\ (10)(2x + 5) = 6x^2 + 11x - 10\) are \( (-2.5, 4)\). So your answer will be C @ashlyn000